OpenStudy (inexilewetrust):

Answer check? Determine velocity of a 2.5 kg object if its momentum is 8520 g * m/s. My Answer: 3408 m/s

OpenStudy (shamim):

Given m=2.5kg p=8520gm.m./s=8.52kg.m/s v=?

OpenStudy (shamim):

Use p=mv v=p/m=?

OpenStudy (inexilewetrust):

Thanks!

OpenStudy (shamim):

U r most welcome!!